OpenStudy (keke):

simplify; 3√(45)-2√(20)

OpenStudy (anonymous):

\[5\sqrt{5}\]

OpenStudy (anonymous):

\[3*3\sqrt{5}-2*2\sqrt{5}=(9-4)\sqrt{5}=5\sqrt{5}\]

OpenStudy (anonymous):

ANAND SIMPLIFY DOES NOT MEAN SOLVING : SAID BY KEKE